# DeID-GPT: Zero-shot Medical Text De-Identification by GPT-4

### v2

@@ I. Introduction
The widespread digitization of medical data has revolutionized healthcare by enabling the easy and efficient sharing of patient information 1 2. The corresponding electronic health record (EHR) systems offer a promising repository of data that can be utilized to expedite the implementation of data-driven solutions and research. At the same time, it has also raised concerns regarding the privacy and security of sensitive medical information 3 4. For example, clinical notes, including physician consultation, nursing assessments, discharge reports, procedure and operative reports, and radiology/pathology reports, are typically archived in a free-text format that frequently incorporates identifiable or confidential patient information. As such, unauthorized access to this information can pose significant risks to patients’ confidential information and privacy 5 6. Patient privacy is always the top concern when sharing, uploading, and processing health information. The U.S. Health Insurance Portability and Accountability Act (HIPAA) mandates the removal of 18 categories of re-identifying information from medical records before their dissemination to preserve the confidentiality of patients 7. It has been an urgent and essential topic for researchers to study potential ways to mitigate related concerns to apply data masking techniques to conceal sensitive data from unauthorized access 8 9.
@@
Recently, large language models (LLM), such as OpenAI’s ChatGPT and GPT-4, have shown tremendous potential in analyzing and processing textual data, thus providing opportunities for downstream tasks in medical data analysis 10; 11; 12; 13; 14; 15; 16; 17. LLMs such as ChatGPT can engage in conversation and generate contextual responses in a more naturalistic manner 18. The latest released GPT-4 can generate, edit, and collaborate with users on creative and technical writing tasks19; 20; 21; 22, as well as interpret and generate images 23; 24. New interesting findings have revealed that LLMs demonstrate a notable aptitude for in-context zero-shot and few-shot learning 25; 26; 27; 28; 29. This discovery boosts the development of the ‘prompt engineering’ technique, which involves providing the LLMs with a brief contextual cue to aid in addressing the given tasks. Due to their ability to generate coherent and contextually appropriate responses, LLMs have been used to produce clinical notes and radiology reports 30; 31; 32. Surprisingly, ChatGPT and GPT-4 also reveal their potential to mask sensitive or private information in medical data while preserving the overall structure and meaning of the text. Given the promising advantages of the LLMs, the objective of the work is to streamline research on clinical notes, particularly narrative free-text notes such as physician and nursing notes, by de-identifying large-scale patients’ sensitive information accurately and efficiently, without manual intervention before medical data sharing and re-using to meet the HIPAA requirements. 33.
@@
The LLM-based methods have several potential advantages in privacy protection. First of all, LLMs have better accuracy in identifying confidential information. Large language models can leverage their ability to learn from vast amounts of data to identify patterns and relationships between words, phrases, and other elements of the text. This makes them well-suited for de-identification tasks, where the goal is to remove identifying information while preserving the meaning of the text. For example, a large language model can identify patterns in names, addresses, phone numbers, and other sensitive information and automatically remove them from text data 34. Second, LLMs can process text data at high speed, making them an efficient tool for de-identification tasks. This is especially important when dealing with large datasets, where manual de-identification can be time-consuming and error-prone. By using a large language model, organizations, including healthcare systems, can quickly process large amounts of text data, de-identify medical records, and reduce the risk of exposing sensitive information while also complying with data privacy regulations such as HIPAA35 36. Third, LLMs can be trained on different types of text data and can learn from a broad range of text data, which enables them to generalize to new, unseen data more effectively, allowing them to adapt to different de-identification tasks and use cases. By training the model on a diverse range of data, it can learn to recognize and remove a wide variety of identifying information. For example, a model trained on medical records can be used to de-identify patient information, while a model trained on financial documents can be used to de-identify credit card numbers and other financial information 37.
@@ II. Related Work
Recently, LLMs in the natural language processing (NLP) domain have been gaining significant attention from the general public. One notable example is ChatGPT, a large language model developed by OpenAI that has been a game-changer in the field of conversational AI, allowing for more natural and human-like interactions between machines and humans. The impressive achievements of ChatGPT have fueled the belief that we are entering a new era of AI, with potentially significant impacts on our society as a whole. By reflecting on the development history of language models, we can gain valuable insights into the evolution of NLP and the ongoing challenges and opportunities in the field. It is crucial to continue to innovate and push the boundaries of what is possible, while also being mindful of the ethical considerations that come with the increasing sophistication of these models. Traditional methods for generating sentences involved learning word distribution using N-gram language modeling 38, followed by a search for the best sequence. However, this method was limited in its ability to adapt to long sentences. To address this problem, recurrent neural networks (RNNs) 39 were introduced for language modeling tasks, allowing for the modeling of relatively long sentences. This was followed by the development of Long Short-Term Memory (LSTM) 40 and Gated Recurrent Unit (GRU) 41, which leveraged gating mechanisms to control memory during training. These approaches could attend to around 200 tokens in a sample 42, representing significant improvement over N-gram language models.
@@
In 2017, the Transformer model was introduced by Vaswani et al. for NLP tasks and has since become the backbone of many LLMs 43; 44; 45; 46; 47; 48. The Transformer model was developed to overcome the limitations of traditional models such as RNNs in handling variable-length sequences and context awareness. It consists of an encoder and an decoder, with the former taking in the input sequence and generating hidden representations, and the latter taking in the hidden representations and producing an output sequence. Each layer of the encoder and decoder contains a multi-head attention and a feed-forward neural network. The multi-head attention is the central component of the Transformer and allows the model to assign different weights to tokens based on their relevance, which helps to improve its performance in handling long-term dependencies in a wide range of NLP tasks. Another advantage of the Transformer is its highly parallelizable architecture, which enables large-scale pre-training and makes it adaptable to various downstream tasks while allowing data to trump inductive biases 49.
@@
Since its introduction, the Transformer architecture has become the dominant choice in NLP due to its parallelism and learning capabilities. Transformer-based pre-trained language models can be broadly classified into two types based on their training tasks: autoregressive language modeling (decoders) and masked language modeling (encoders) 50. Masked language modeling, such as BERT 43 and RoBERTa 44, involves predicting the probability of a masked token given contextual information. In contrast, autoregressive language modeling, such as the GPT family 45; 46; 47 and OPT 48, is focused on modeling the probability of the next token given the previous tokens, and is better suited for generative tasks. The GPT model, a transformer-based autoregressive decoder model that uses self-attention mechanisms to process all words in a sequence simultaneously, is one of the most notable examples of autoregressive language modeling. GPT is trained on a next word prediction task based on previous tokens, allowing it to generate coherent text. Subsequently, GPT-2 47 and GPT-3 45 continue to use the autoregressive left-to-right training method, while scaling up the number of model parameters and incorporating diverse datasets beyond basic web text, achieving state-of-the-art results on numerous NLP tasks. In addition to the two types of models based on the encoder or decoder architecture of Transformer, there are several LLMs that employ a complete encoder-decoder structure. Text-to-Text Transfer Transformer (T5) 51 and BART 52 are two examples of such models. T5 is one of the primary encoder-decoder methods that utilizes a ”text-to-text” approach. This means that both input and output data are transformed into a standardized text format, enabling T5 to be trained on various NLP tasks, including machine translation, question-answering, and summarization, using the same model architecture. Another frequently used encoder-decoder method is BART, which combines the bidirectional encoder from BERT and the autoregressive decoder from GPT. BART takes advantage of the bidirectional modeling abilities of the encoder while retaining the autoregressive properties for generative tasks.
@@
Recently, a number of very large language models have emerged, including Megatron-Turing Natural Language Generation (MT-NLG), a monolithic transformer English language model with 530 billion parameters that outperforms prior state-of-the-art models in zero-, one-, and few-shot settings 53. Pathways Language Model (PaLM), a dense decoder-only Transformer model, has 540 billion parameters and is trained with the pathways system. It achieved state-of-the-art few-shot performance across most tasks, with significant margins in many cases 54. Additionally, BigScience Large Open-science Open-access Multilingual Language Model (BLOOM) is another transformer-based large language model, which is trained on approximately 176 billion parameters 55. Furthermore, Jurassic-1 models come in two sizes, with the Jumbo version (178 billion parameters) being the largest and most sophisticated language model ever released for general use by developers 56.
@@
Recently, with the development of LLMs like GPT-3 and GPT-4, which are pretrained on massive datasets and capable of capturing a wide range of tasks and knowledge, zero-shot and few-shot in-context learning have become feasible, pushing LLMs into real-world applications. This is because these models leverage prior knowledge obtained from pretraining on diverse tasks, allowing them to quickly adapt to new tasks without the need for collecting labeled data for extensive fine-tuning, which can be particularly challenging in medical fields where labeled data are limited or not available at all 57.
Specifically, for LLMs in NLP, zero-shot and few-shot in-context learning refer to the model’s ability to understand and perform a new task by simply providing a few examples of the desired input-output pairs  11 within the prompt, or even just the task instructions without any examples. The prompts help the model grasp the structures and patterns of the task, while zero-shot and few-shot in-context learning behave similarly to explicit fine-tuning at the prediction level, the representation level and the attention behavior level, enabling it to generalize and perform the new task even better without further training or fine-tuning 58, and reducing the possibility of overfitting on downstream labeled training data. While no fine-tuning is needed for these LLMs, the trade-offs include increased computational costs during inference and the potential need for expert knowledge to craft effective prompts with examples.
@@
LLMs are a promising tool in domains where humans and AI work together to create software-reliant systems more quickly and reliably 59. However, the process of collecting and labeling responses for training or fine-tuning NLP models is time-consuming and costly. Recent studies suggest that large-scale pre-trained language models (PLMs) can be adapted to downstream tasks without fine-tuning by using prompts 60.
A prompt is a set of instructions that customizes an LLM’s capabilities and influences subsequent interactions and outputs. Prompts can do more than just filter information or dictate output types; they can be engineered to create entirely new interaction paradigms, like generating quizzes or simulating a Linux terminal window 61. Additionally, prompts have the potential to suggest other prompts for self-adaptation, making them a valuable tool in NLP 61. In general, prompt engineering leads to a new paradigm in NLP. The advanced capabilities of prompts highlight the importance of engineering them to provide values beyond simple text or code generation. However, finding the most appropriate prompt presents a new challenge. Currently, prompts can be either created manually 62; 63 or learned automatically 60; 64. While automatically learned prompts can yield better performance in some tasks, they are often not human-readable. Therefore, in domains where interpretability is essential, such as medical domains, manually created prompts are more commonly used.
Although prompt engineering is still at its early stage, some valuable insights into effective prompt patterns have been proposed 61. For example, researchers in 61 compared prompt patterns to well-known software patterns, as both offer reusable solutions to common issues in a specific context, such as output generation and interaction when working with LLMs. They described 16 prompt patterns that have proven effective in improving the outputs of LLM-based conversations. These patterns have provided a great deal of inspiration, and as the field advances, it is likely that even more effective prompt patterns will be discovered.
@@
Significant progresses have been made in filtering out confidential information from medical data, but they are usually time-consuming and far from satisfactory in practical use 65 66. The existing de-identification methods for medical text data are mainly manual, rule-based, and learning-based. Figure 2 shows the brief development of de-identification methods.
@@
Manual Methods The manual de-identification of EHRs involves human annotators who read and remove sensitive information. However, this approach is both costly and time-consuming. The existing report shows that the cost of using human annotators is \$50 per hour, and they can read around 20,000 words per hour 67. For a dataset containing 100 million words, the cost would be a staggering number of \$250,000. Additionally, the recall value for manual de-identification ranges from 63% to 94%, depending on the annotators 68. Besides, the manual de-identification process is prone to errors, making it an unreliable and inefficient approach.
@@
Rule-based Methods Considering the limitations of manual de-identification, automated approaches have emerged as a logical advancement for automatically de-identifying EHRs. These automated de-identification approaches are predominantly rule-based. Rule-based methods mainly depend on pre-defined word patterns with regular expressions and look-up searching dictionaries. For example, pseudonymisation 69 aims to de-identify clinical data by either removing entire sentences that contain sensitive information or by substituting sensitive words with realistic alternatives, which may undermine the performance of the model 70. A hybrid model was proposed that utilizes machine learning techniques in conjunction with keyword-based and rule-based approaches, incorporating a diverse array of linguistic features, task-specific features, and regular expression template patterns to effectively address the complexities inherent in personal health information categorization 71. While these systems are relatively simple to create, they lack generalizability due to the need for fine-tuning the rules for each dataset, and they fail to consider the contextual nuances of words.
@@
Learning-based Methods Learning-based methods built scalable pipelines for de-identification using machine learning and deep learning methods 72 73, however, this usually results in unstable performance when the real data has a domain shift from the training data 74 75. Though deep-learning-based methods with advanced techniques from the NLP community have been proposed to improve de-identification ability on cross-institute datasets, these methods are only trained and tested on rather small datasets and depend heavily on pre-preprocessing pipelines, such as pre-processing clinical notes 75. During the COVID-19 pandemic, cross-lingual transfer learning was applied to de-identify medical records written in a low-resource language using the models built with the data written in high resources languages 76. Improved named entity recognition (NER) model was proposed that incorporated a multi-faceted input embedding layer, consisting of standard word embedding, context-based word embedding, character-level word embedding via a convolutional neural network (CNN), external knowledge sources, and one-hot vector word features 77.
@@
Healthcare: A prominent application prospect of LLMs is LLM-powered healthcare. Clinical practices generate exceedingly large amounts of texts on a daily basis. It is beneficial to unlock the potential of clinical text data through state-of-the-art tools and methods. The inception of the Transformer  78 has led to the rise of pre-trained language models in healthcare  79; 80. For example, the ClinicalRadioBERT is a specialized language model for radiation oncology  81. Other applications include radiology report summarization  82, mental disorder detection  83, COVID-19 research summarization  84, and clinical information extraction  85. However, the scale of existing language models is not comparable to recent developments such as ChatGPT and GPT-4. Indeed, larger models have learned from more materials and typically have more capacity to handle diverse and complex tasks. Therefore, we foresee broad adoption of ChatGPT and GPT-4 in a wide range of tasks, ranging from medical triage and question answering for patients, to knowledge-guided physicians assistance tools and fully-automated information extraction and collection. In addition, the recent popularity of LLMs is also pushing the boundary of language models, because the widespread success might break barriers that have prevented in-depth adoption of LLM-based methods  86; 87. The next generation of ChatGPT will be powered by GPT-4, which is a considerable upgrade from GPT-3.5, and we believe this will bring more innovations to healthcare NLP. Finally, it should be noted that ChatGPT is trained with RLHF. This process incorporates human preference, human values and guardrails (e.g., the exclusion of toxic generation results) into this LLM. This capability distinguishes ChatGPT from other LLMs, especially since it is extremely valuable for LLMs to understand human values and patient-centered guidelines.

### v3

@@ I. Introduction
The widespread digitization of medical data has revolutionized healthcare by enabling the easy and efficient sharing of patient information Roberto23 mcdowell2022oral. The corresponding electronic health record (EHR) systems offer a promising repository of data that can be utilized to expedite the implementation of data-driven solutions and research. At the same time, it has also raised concerns regarding the privacy and security of sensitive medical information Tang23 Paul23. For example, clinical notes, including physician consultation, nursing assessments, discharge reports, procedure and operative reports, and radiology/pathology reports, are typically archived in a free-text format that frequently incorporates identifiable or confidential patient information. As such, unauthorized access to this information can pose significant risks to patients’ confidential information and privacy urbain2022natural tayefi2021challenges. Patient privacy is always the top concern when sharing, uploading, and processing health information. The U.S. Health Insurance Portability and Accountability Act (HIPAA) mandates the removal of 18 categories of re-identifying information from medical records before their dissemination to preserve the confidentiality of patients Tanbir20. It has been an urgent and essential topic for researchers to study potential ways to mitigate related concerns to apply data masking techniques to conceal sensitive data from unauthorized access Nandita21 Asokan19.
@@
Recently, large language models (LLM), such as OpenAI’s ChatGPT and GPT-4, have shown tremendous potential in analyzing and processing textual data, thus providing opportunities for downstream tasks in medical data analysis thoppilan2022lamda; dai2023chataug; liu2023artificial; li2023artificial1; liu2023radonc; liu2023evaluating; liu2023summary; holmes2023evaluating. LLMs such as ChatGPT can engage in conversation and generate contextual responses in a more naturalistic manner Long22. The latest released GPT-4 can generate, edit, and collaborate with users on creative and technical writing taskszhong2023chatradio; liu2023tailoring; liu2023radiology; liu2023radiology1, as well as interpret and generate images liu2023holistic; yan2023multimodal. New interesting findings have revealed that LLMs demonstrate a notable aptitude for in-context zero-shot and few-shot learning Takeshi22; Monica22; Hongjin22; shi2023mededit; guan2023cohortgpt. This discovery boosts the development of the ‘prompt engineering’ technique, which involves providing the LLMs with a brief contextual cue to aid in addressing the given tasks. Due to their ability to generate coherent and contextually appropriate responses, LLMs have been used to produce clinical notes and radiology reports jeblick2022chatgpt; Som23; ma2023impressiongpt. Surprisingly, ChatGPT and GPT-4 also reveal their potential to mask sensitive or private information in medical data while preserving the overall structure and meaning of the text. Given the promising advantages of the LLMs, the objective of the work is to streamline research on clinical notes, particularly narrative free-text notes such as physician and nursing notes, by de-identifying large-scale patients’ sensitive information accurately and efficiently, without manual intervention before medical data sharing and re-using to meet the HIPAA requirements. Carmel22.
@@
The LLM-based methods have several potential advantages in privacy protection. First of all, LLMs have better accuracy in identifying confidential information. Large language models can leverage their ability to learn from vast amounts of data to identify patterns and relationships between words, phrases, and other elements of the text. This makes them well-suited for de-identification tasks, where the goal is to remove identifying information while preserving the meaning of the text. For example, a large language model can identify patterns in names, addresses, phone numbers, and other sensitive information and automatically remove them from text data Chengwei23. Second, LLMs can process text data at high speed, making them an efficient tool for de-identification tasks. This is especially important when dealing with large datasets, where manual de-identification can be time-consuming and error-prone. By using a large language model, organizations, including healthcare systems, can quickly process large amounts of text data, de-identify medical records, and reduce the risk of exposing sensitive information while also complying with data privacy regulations such as HIPAAFares23 Maad23. Third, LLMs can be trained on different types of text data and can learn from a broad range of text data, which enables them to generalize to new, unseen data more effectively, allowing them to adapt to different de-identification tasks and use cases. By training the model on a diverse range of data, it can learn to recognize and remove a wide variety of identifying information. For example, a model trained on medical records can be used to de-identify patient information, while a model trained on financial documents can be used to de-identify credit card numbers and other financial information Xiang23.
@@ II. Related Work
Recently, LLMs in the natural language processing (NLP) domain have been gaining significant attention from the general public. One notable example is ChatGPT, a large language model developed by OpenAI that has been a game-changer in the field of conversational AI, allowing for more natural and human-like interactions between machines and humans. The impressive achievements of ChatGPT have fueled the belief that we are entering a new era of AI, with potentially significant impacts on our society as a whole. By reflecting on the development history of language models, we can gain valuable insights into the evolution of NLP and the ongoing challenges and opportunities in the field. It is crucial to continue to innovate and push the boundaries of what is possible, while also being mindful of the ethical considerations that come with the increasing sophistication of these models. Traditional methods for generating sentences involved learning word distribution using N-gram language modeling bengio2000neural, followed by a search for the best sequence. However, this method was limited in its ability to adapt to long sentences. To address this problem, recurrent neural networks (RNNs) mikolov2010recurrent were introduced for language modeling tasks, allowing for the modeling of relatively long sentences. This was followed by the development of Long Short-Term Memory (LSTM) graves2012long and Gated Recurrent Unit (GRU) dey2017gate, which leveraged gating mechanisms to control memory during training. These approaches could attend to around 200 tokens in a sample khandelwal2018sharp, representing significant improvement over N-gram language models.
@@
In 2017, the Transformer model was introduced by Vaswani et al. for NLP tasks and has since become the backbone of many LLMs devlin2018bert; liu2019roberta; brown2020language; radford2018improving; radford2019language; zhang2022opt. The Transformer model was developed to overcome the limitations of traditional models such as RNNs in handling variable-length sequences and context awareness. It consists of an encoder and an decoder, with the former taking in the input sequence and generating hidden representations, and the latter taking in the hidden representations and producing an output sequence. Each layer of the encoder and decoder contains a multi-head attention and a feed-forward neural network. The multi-head attention is the central component of the Transformer and allows the model to assign different weights to tokens based on their relevance, which helps to improve its performance in handling long-term dependencies in a wide range of NLP tasks. Another advantage of the Transformer is its highly parallelizable architecture, which enables large-scale pre-training and makes it adaptable to various downstream tasks while allowing data to trump inductive biases elhage2021mathematical.
@@
Since its introduction, the Transformer architecture has become the dominant choice in NLP due to its parallelism and learning capabilities. Transformer-based pre-trained language models can be broadly classified into two types based on their training tasks: autoregressive language modeling (decoders) and masked language modeling (encoders) qiu2020pre. Masked language modeling, such as BERT devlin2018bert and RoBERTa liu2019roberta, involves predicting the probability of a masked token given contextual information. In contrast, autoregressive language modeling, such as the GPT family brown2020language; radford2018improving; radford2019language and OPT zhang2022opt, is focused on modeling the probability of the next token given the previous tokens, and is better suited for generative tasks. The GPT model, a transformer-based autoregressive decoder model that uses self-attention mechanisms to process all words in a sequence simultaneously, is one of the most notable examples of autoregressive language modeling. GPT is trained on a next word prediction task based on previous tokens, allowing it to generate coherent text. Subsequently, GPT-2 radford2019language and GPT-3 brown2020language continue to use the autoregressive left-to-right training method, while scaling up the number of model parameters and incorporating diverse datasets beyond basic web text, achieving state-of-the-art results on numerous NLP tasks. In addition to the two types of models based on the encoder or decoder architecture of Transformer, there are several LLMs that employ a complete encoder-decoder structure. Text-to-Text Transfer Transformer (T5) raffel2020exploring and BART lewis2019bart are two examples of such models. T5 is one of the primary encoder-decoder methods that utilizes a ”text-to-text” approach. This means that both input and output data are transformed into a standardized text format, enabling T5 to be trained on various NLP tasks, including machine translation, question-answering, and summarization, using the same model architecture. Another frequently used encoder-decoder method is BART, which combines the bidirectional encoder from BERT and the autoregressive decoder from GPT. BART takes advantage of the bidirectional modeling abilities of the encoder while retaining the autoregressive properties for generative tasks.
@@
Recently, a number of very large language models have emerged, including Megatron-Turing Natural Language Generation (MT-NLG), a monolithic transformer English language model with 530 billion parameters that outperforms prior state-of-the-art models in zero-, one-, and few-shot settings Smith2022. Pathways Language Model (PaLM), a dense decoder-only Transformer model, has 540 billion parameters and is trained with the pathways system. It achieved state-of-the-art few-shot performance across most tasks, with significant margins in many cases Chowdhery2022PaLM. Additionally, BigScience Large Open-science Open-access Multilingual Language Model (BLOOM) is another transformer-based large language model, which is trained on approximately 176 billion parameters Laurencon2023. Furthermore, Jurassic-1 models come in two sizes, with the Jumbo version (178 billion parameters) being the largest and most sophisticated language model ever released for general use by developers Lieber2021.
@@
Recently, with the development of LLMs like GPT-3 and GPT-4, which are pretrained on massive datasets and capable of capturing a wide range of tasks and knowledge, zero-shot and few-shot in-context learning have become feasible, pushing LLMs into real-world applications. This is because these models leverage prior knowledge obtained from pretraining on diverse tasks, allowing them to quickly adapt to new tasks without the need for collecting labeled data for extensive fine-tuning, which can be particularly challenging in medical fields where labeled data are limited or not available at all zhang2023biomedgpt.
Specifically, for LLMs in NLP, zero-shot and few-shot in-context learning refer to the model’s ability to understand and perform a new task by simply providing a few examples of the desired input-output pairs  dai2023chataug within the prompt, or even just the task instructions without any examples. The prompts help the model grasp the structures and patterns of the task, while zero-shot and few-shot in-context learning behave similarly to explicit fine-tuning at the prediction level, the representation level and the attention behavior level, enabling it to generalize and perform the new task even better without further training or fine-tuning dai2022can, and reducing the possibility of overfitting on downstream labeled training data. While no fine-tuning is needed for these LLMs, the trade-offs include increased computational costs during inference and the potential need for expert knowledge to craft effective prompts with examples.
@@
LLMs are a promising tool in domains where humans and AI work together to create software-reliant systems more quickly and reliably carleton2022architecting. However, the process of collecting and labeling responses for training or fine-tuning NLP models is time-consuming and costly. Recent studies suggest that large-scale pre-trained language models (PLMs) can be adapted to downstream tasks without fine-tuning by using prompts liu2023pre.
A prompt is a set of instructions that customizes an LLM’s capabilities and influences subsequent interactions and outputs. Prompts can do more than just filter information or dictate output types; they can be engineered to create entirely new interaction paradigms, like generating quizzes or simulating a Linux terminal window white2023prompt. Additionally, prompts have the potential to suggest other prompts for self-adaptation, making them a valuable tool in NLP white2023prompt. In general, prompt engineering leads to a new paradigm in NLP. The advanced capabilities of prompts highlight the importance of engineering them to provide values beyond simple text or code generation. However, finding the most appropriate prompt presents a new challenge. Currently, prompts can be either created manually liu2021gpt; schick2020s or learned automatically liu2023pre; shin2020autoprompt. While automatically learned prompts can yield better performance in some tasks, they are often not human-readable. Therefore, in domains where interpretability is essential, such as medical domains, manually created prompts are more commonly used.
Although prompt engineering is still at its early stage, some valuable insights into effective prompt patterns have been proposed white2023prompt. For example, researchers in white2023prompt compared prompt patterns to well-known software patterns, as both offer reusable solutions to common issues in a specific context, such as output generation and interaction when working with LLMs. They described 16 prompt patterns that have proven effective in improving the outputs of LLM-based conversations. These patterns have provided a great deal of inspiration, and as the field advances, it is likely that even more effective prompt patterns will be discovered.
@@
Significant progresses have been made in filtering out confidential information from medical data, but they are usually time-consuming and far from satisfactory in practical use Thomas22 Iyadh22. The existing de-identification methods for medical text data are mainly manual, rule-based, and learning-based. Figure 2 shows the brief development of de-identification methods.
@@
Manual Methods The manual de-identification of EHRs involves human annotators who read and remove sensitive information. However, this approach is both costly and time-consuming. The existing report shows that the cost of using human annotators is \$50 per hour, and they can read around 20,000 words per hour Cliffford05. For a dataset containing 100 million words, the cost would be a staggering number of \$250,000. Additionally, the recall value for manual de-identification ranges from 63% to 94%, depending on the annotators Ishna08. Besides, the manual de-identification process is prone to errors, making it an unreliable and inefficient approach.
@@
Rule-based Methods Considering the limitations of manual de-identification, automated approaches have emerged as a logical advancement for automatically de-identifying EHRs. These automated de-identification approaches are predominantly rule-based. Rule-based methods mainly depend on pre-defined word patterns with regular expressions and look-up searching dictionaries. For example, pseudonymisation Hercules19 aims to de-identify clinical data by either removing entire sentences that contain sensitive information or by substituting sensitive words with realistic alternatives, which may undermine the performance of the model Hanna20. A hybrid model was proposed that utilizes machine learning techniques in conjunction with keyword-based and rule-based approaches, incorporating a diverse array of linguistic features, task-specific features, and regular expression template patterns to effectively address the complexities inherent in personal health information categorization Hui15Automatric. While these systems are relatively simple to create, they lack generalizability due to the need for fine-tuning the rules for each dataset, and they fail to consider the contextual nuances of words.
@@
Learning-based Methods Learning-based methods built scalable pipelines for de-identification using machine learning and deep learning methods Joffrey20 Junhak22, however, this usually results in unstable performance when the real data has a domain shift from the training data Urbain22 Xi19. Though deep-learning-based methods with advanced techniques from the NLP community have been proposed to improve de-identification ability on cross-institute datasets, these methods are only trained and tested on rather small datasets and depend heavily on pre-preprocessing pipelines, such as pre-processing clinical notes Xi19. During the COVID-19 pandemic, cross-lingual transfer learning was applied to de-identify medical records written in a low-resource language using the models built with the data written in high resources languages catelli2020crosslingual. Improved named entity recognition (NER) model was proposed that incorporated a multi-faceted input embedding layer, consisting of standard word embedding, context-based word embedding, character-level word embedding via a convolutional neural network (CNN), external knowledge sources, and one-hot vector word features syed2022deidner.
@@
Healthcare: A prominent application prospect of LLMs is LLM-powered healthcare. Clinical practices generate exceedingly large amounts of texts on a daily basis. It is beneficial to unlock the potential of clinical text data through state-of-the-art tools and methods. The inception of the Transformer  vaswani2017attention has led to the rise of pre-trained language models in healthcare  wu2022survey; liu2022survey. For example, the ClinicalRadioBERT is a specialized language model for radiation oncology  rezayi2022clinicalradiobert. Other applications include radiology report summarization  cai2021chestxraybert, mental disorder detection  ji2021mentalbert, COVID-19 research summarization  cai2022covidsum, and clinical information extraction  agrawal2022large. However, the scale of existing language models is not comparable to recent developments such as ChatGPT and GPT-4. Indeed, larger models have learned from more materials and typically have more capacity to handle diverse and complex tasks. Therefore, we foresee broad adoption of ChatGPT and GPT-4 in a wide range of tasks, ranging from medical triage and question answering for patients, to knowledge-guided physicians assistance tools and fully-automated information extraction and collection. In addition, the recent popularity of LLMs is also pushing the boundary of language models, because the widespread success might break barriers that have prevented in-depth adoption of LLM-based methods  abdullah2023chatgpt; haleem2023era. The next generation of ChatGPT will be powered by GPT-4, which is a considerable upgrade from GPT-3.5, and we believe this will bring more innovations to healthcare NLP. Finally, it should be noted that ChatGPT is trained with RLHF. This process incorporates human preference, human values and guardrails (e.g., the exclusion of toxic generation results) into this LLM. This capability distinguishes ChatGPT from other LLMs, especially since it is extremely valuable for LLMs to understand human values and patient-centered guidelines.